OpenStudy (anonymous):

polpak help me : u said that log(1+2+3) = log1+ log2+ log3 only in this case . my question is : can this be proved ?

OpenStudy (anonymous):

what is log 6?

OpenStudy (anonymous):

I'l prove it with a calculator =P

OpenStudy (anonymous):

polpak said me that log(1+2+3) = log 6 or log1+log2+log3 , but only for the case of 1 , 2 and 3 he said that i am asking , that can it be proved

OpenStudy (anonymous):

if its only for this case then calculation suffices as proof

OpenStudy (anonymous):

log6=log(1*2*3) =log1+log2+log3

OpenStudy (anonymous):

that's the proof

OpenStudy (anonymous):

1+2+3=1*2*3

OpenStudy (anonymous):

yes log6=log(1+2+3)

OpenStudy (saifoo.khan):

that is the Product rule

OpenStudy (anonymous):

In this case or in any case dealing with log 1 you can do so this because in the product rule expansion you will get log 1 as 0